20220510_DX_Day. LBTI AO, Day 4

15:04 UT Changing ASM configuration to daytime

15:07 UT Fastlink test failed on first program loading attempt, successful the second

15:08 UT applying default flat 565 at +0.17 N (RB temperature 3.9 C)

15:12 UT loaded the /local/aomeas/adsec_calib/flat/20220504_154450/Ref_2C_flat.sav 565 at +0.05 N (set as default)

15:13 UT acquiring the source from the WFS side

15:35 UT source acquisition completed (nominally set to mag 7.5)

15:43 UT first Preset AO (preset shows 1700 Hz, we need to check the AOS table)

15:45 UT the Bayside stage position is incorrect, the source was lost

15:51 UT acquiring a new dark for OCAM2K

15:54 UT trousbleshooting some issue with the OCAM2K

15:55 UT saving some data, to be used later for the troubleshooting of the 'bad frames' OCAM2K issue

15:58 UT second Preset AO (with corrected configuration)

15:59 UT acquired dark is not good

16:01 UT center pupil failed (IDL: invalid incoherent parameter for the reconstructor)

16:08 UT first AO CL failed and sent the mirror position to a max of 92um

16:10 UT second AO CL, using a different reconstructor, seems to work

16:13 UT third AO CL, the pupil seems to be off

16:19 UT 900 Hz 3 lambda/D for the WFS and IIR temporal filtering

16:21 UT the AO CL with previous parameters did not work

16:19 UT 900 Hz 3 lambda/D for the WFS and pureIntegrator temporal filtering

16:22 UT the AO CL with previous parameters did not work (the pupil file was not correct)

16:23 UT re-flattening, force on 565 at +0.02 N (RB temperature 4.4 C)

16:28 UT the AO CL was successful (the problem appeared to have been the wrong pupil file)

16:30 UT 900 Hz 3 lambda/D for the WFS and IIR temporal filtering

16:30 UT the AO CL was successful (you can see that the ASM shows slightly higher forces in the quadrant 6-9 o-clock, could it be a pupil issue?)

16:34 UT it appears that there might be some issues with the AOS configuration, we will troubleshoot this for some time...

16:54 UT discussed with Doug about the set_z0 tests on SX ASM

17:13 UT the WFS was restarted, the AOS table was copied in the correct the directory

17:17 UT resting the shell and cycling the AO Arbitrator and AdSec Arbitrator

17:18 UT sending Preset AO

17:22 UT the problem was that the current configuration RR was using a different AOS table (RR type), now it works through AOS

17:23 UT after making the symbolic link of the used AOS table sending a Preset AO (now it shows the correct pupil and frequency at the AOS Command GUI)

17:25 UT AO CL running with 1" disturbance

17:25 UT gain optimization running

17:27 UT optimization finished, turning on Gopt (temporal filtering IIR with associated reconstructor)

17:27 UT 172720 (2000) without Gopt

17:28 UT 172827, 172849, 172900, 172919 (2000) optimized with Gopt on 1” Dist

17:30 UT zero AO gain 173032, 173044, 173055, 173106 (2000) zero gain 1” Dist

17:35 UT changing configuration to MMC: 565, 367, 340, 115, 484, 628, 126

17:48 UT loaded /local/aomeas/adsec_calib/flat/20220504_154450/Ref_2C_flat.sav 565 at +0.11 N (default)

17:49 UT applied /local/aomeas/adsec_calib/flat/20220422_191142/flattening.sav 565 at -0.04 N (better, set as default) RB temperature 5.0 C

17:52 UT Preset AO

17:56 UT Preset AO was cancelled because it did not have the disturbance (which is unexpected)

17:57 UT Preset AO (RR was enabled) successfully

18:02 UT without Gop 180113, 180125 (2000), 180150 (4000) loop closed with LN gains Dist 1”

18:06 UT 180526, 180543, 180555 GOpt on, Optimized

18:07 UT Zero gain: 180640, 180654, 180706

18:33 UT additional tests are taking place for AO

18:38 UT RB temperature 5.3 C

19:16 UT reverting back to night configuration (the first attempt failed)

19:20 UT configuration change completed

-- GuidoBrusa - 10 May 2022
Topic revision: r2 - 10 May 2022, GuidoBrusa
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback